# On the number of modes of Gaussian kernel density estimators

### v2

@@ Appendix A Additional proofs
$\int_{0}^{\infty}z^{n}e^{-\alpha u^{2}}\,\mathrm{d}u=\frac{1}{2}\Gamma\mathopen{}\mathclose{{}\left(\frac{n+1}{2}}\right)\alpha^{-\frac{n+1}{2}}.$
@@
Hence, using A.1 we compute
$\mathbb{E}G(t)=\int_{-\infty}^{\infty}ze^{-\frac{\beta}{2}z^{2}}\,\mathrm{d}\gamma_{t,1}(z)$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}ze^{-\frac{\beta}{2}z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
@@
$\mathbb{E}G^{\prime}(t)$ $=\int_{-\infty}^{\infty}(1-\beta z^{2})z^{-\frac{\beta}{2}z^{2}}\,\mathrm{d}\gamma_{t,1}(z)$
@@
From these computations, and the remark after A.1, we readily obtain the asymptotics of $\mu_{t}$ as in 2.2 upon multiplying by $\sqrt{n}$.
@@
$z^{2}+\frac{1}{2}(z-t)^{2}=\frac{2\beta+1}{2}u^{2}+\frac{\beta t^{2}}{2\beta+1}\quad\text{where}\quad u=z-\frac{t}{2\beta+1}.$
Hence using A.1 we compute
@@
We check that entries of $\Sigma_{t}$ are asymptotically the corresponding second moments. Indeed, surpressing the dependence on $t$, we have that
  • $\mathbb{E}G^{2}\asymp\beta^{-\frac{5}{2}}(t^{2}+\beta)e^{-\frac{t^{2}}{2}}\gg\beta^{-3}t^{2}e^{-t^{2}}\asymp(\mathbb{E}G)^{2},$
  • $\mathbb{E}GG^{\prime}\asymp-\beta^{-\frac{7}{2}}e^{-\frac{t^{2}}{2}}(\beta^{2}t+\beta t^{3})\gg\beta^{-4}e^{-t^{2}}t(\beta-\beta t^{2})\asymp(\mathbb{E}G)(\mathbb{E}G^{\prime}),$
  • $\mathbb{E}G^{\prime 2}\asymp\beta^{-\frac{9}{2}}e^{-\frac{t^{2}}{2}}(\beta^{4}+\beta^{2}t^{4})\gg\beta^{-5}e^{-t^{2}}\beta^{2}(1+t^{4})\asymp(\mathbb{E}G^{\prime})^{2},$
where we bound $e^{-\frac{t^{2}}{2}}\leqslant 1$. From these computations, and the remark after A.1, we readily obtain the asymptotics of $\Sigma_{t}$ as indicated in 2.2.∎
A.2 Proof of 3.1
@@
In this section, we prove 3.1 on third moments of $Y=\Sigma_{t}^{-1/2}\mathopen{}\mathclose{{}\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}\right)$. To upper bound, we do not need to track the leading coefficients to ensure that they do not vanish when we combine applications of A.1. Recalling $\Sigma_{t}^{-1}$ from the proof of Lemma 2.3, we upper bound asymptotically via Hölder’s inequality:
$\eta_{3}$ $=\max_{k}\mathopen{}\mathclose{{}\left|Y^{(k,3-k)}}\right|$
$\leqslant\mathbb{E}\|Y\|^{3}$
@@
$\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left\|\mathopen{}\mathclose{{}\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}\right)^{\intercal}\Sigma_{t}^{-1}\mathopen{}\mathclose{{}\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}\right)}\right\|^{\frac{3}{2}}$
@@
$\lesssim\beta^{\frac{3}{4}}e^{\frac{3t^{2}}{4}}\mathbb{E}\mathopen{}\mathclose{{}\left|\beta(G-\mathbb{E}G)^{2}+2t(G-\mathbb{E}G)(G^{\prime}-\mathbb{E}G^{\prime})+(G^{\prime}-\mathbb{E}G^{\prime})^{2}}\right|^{\frac{3}{2}}$
@@
$\lesssim\beta^{\frac{3}{4}}e^{\frac{3t^{2}}{4}}\mathopen{}\mathclose{{}\left(\beta^{\frac{3}{2}}\mathbb{E}|G|^{3}+\mathbb{E}|G^{\prime}|^{3}}\right)$
@@
$\lesssim\beta^{\frac{3}{4}}e^{\frac{3t^{2}}{4}}\int_{-\infty}^{\infty}\mathopen{}\mathclose{{}\left(\beta^{\frac{3}{2}}|z|^{3}+|1-\beta z^{2}|^{3}}\right)e^{-\frac{3\beta}{2}z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
$\lesssim\beta^{\frac{3}{4}}e^{\frac{t^{2}}{4}}\int_{0}^{\infty}h\mathopen{}\mathclose{{}\left(u+\frac{t}{3\beta+1}}\right)e^{-\frac{3\beta+1}{2}u^{2}}\,\mathrm{d}u,$
where we factor out $e^{-\frac{3\beta}{2(3\beta+1)}t^{2}}\asymp e^{-\frac{t^{2}}{2}}$ of the integral by substituting
$h(z)=\beta^{\frac{3}{2}}z^{3}+(1+\beta z^{2})^{3}\quad\text{and}\quad u=z-\frac{t}{3\beta+1}.$
@@
By linearity of integration and A.1, each monomial $u^{n}$ integrates to $O(\beta^{-(n+1)/2})$. By monotonicity of $h$ on $\mathbb{R}_{\geqslant 0}$ and since $t\in T$, for some constant $C>0$,
$\eta_{3}\lesssim\beta^{\frac{3}{4}}e^{\frac{t^{2}}{4}}\int_{0}^{\infty}h\mathopen{}\mathclose{{}\left(u+\frac{t}{3\beta+1}}\right)e^{-\frac{3\beta+1}{2}u^{2}}\,\mathrm{d}u$ $\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}h\mathopen{}\mathclose{{}\left(C\beta^{-\frac{1}{2}}+\frac{t}{3\beta+1}}\right)$
$\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}h\mathopen{}\mathclose{{}\left(2C\beta^{-\frac{1}{2}}}\right)$
$\lesssim\beta^{\frac{1}{4}}e^{\frac{t^{2}}{4}}$
@@
upon noting $w\mapsto h(w/\sqrt{\beta})$ has constant coefficients. This proves 3.1.∎
A.3 Proof of Lemma 3.4
@@
Fix $n,\beta$ sufficiently large. For a multi-index $\alpha$, define
$h(\mathbf{x})$ $=\mathbf{x}^{\alpha}\mathopen{}\mathclose{{}\left(q_{t}-\varphi-n^{-\frac{1}{2}}\psi}\right)(\mathbf{x}),$
@@
$\mathscr{F}h(\mathbf{z})$ $=\partial^{\alpha}\mathopen{}\mathclose{{}\left(\mathscr{F}(q_{t})-\frac{\varphi}{6\sqrt{n}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}}\right)(\mathbf{z}),$
where
$(\mathscr{F}h)(\mathbf{z})=\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\langle\mathbf{x},\mathbf{z}\rangle}h(\mathbf{x})\,\mathrm{d}\mathbf{x}$
@@
denotes the Fourier transform of $h$. By Fourier inversion, it suffices to show that for any multi-index $\alpha$ with order $|\alpha|\leqslant 3$,
@@
$\mathopen{}\mathclose{{}\left|h(\mathbf{x})}\right|=\mathopen{}\mathclose{{}\left|\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\langle\mathbf{z},\mathbf{x}\rangle}\mathscr{F}{h}(\mathbf{z})\,\mathrm{d}\mathbf{z}}\right|\lesssim\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\mathscr{F}{h}(\mathbf{z})}\right|\,\mathrm{d}\mathbf{z}$ (A.1)
@@
is $O(\exp(-\omega(\beta)/2))$. We apply [BR10, Theorem 9.12]—which is not asymptotic and has explicit constants—so we may use it even though $q_{t}$ depends on $n$. Upon verifying the conditions on $Y$ via 2.2 and 3.1, we have that
@@
$\mathopen{}\mathclose{{}\left|\mathscr{F}{h}(\mathbf{z})}\right|\lesssim n^{-\frac{1}{2}}\eta_{3}^{\frac{1}{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{4}}$
@@
provided $\|\mathbf{z}\|\leqslant a\sqrt{n}$ for some $a\asymp\eta_{3}^{-1/2}$. By 3.1, we have that
@@
$\int_{\|\mathbf{z}\|\leqslant a\sqrt{n}}\mathopen{}\mathclose{{}\left|\mathscr{F}{h}(\mathbf{z})}\right|\,\mathrm{d}\mathbf{z}\lesssim n^{-\frac{1}{2}}\eta_{3}^{\frac{1}{2}}\int_{\mathbb{R}^{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{4}}\,\mathrm{d}\mathbf{z}\lesssim e^{-\frac{\omega(\beta)}{2}}.$ (A.2)
Recall that $q_{t}$ is the density of $n^{-1/2}\sum_{i=1}^{n}Y_{i}$, and let $f$ denote the density of $Y$. Now, we proceed as the proof of [BR10, Theorem 19.2]. As $p_{t}$ is bounded, so is $f$, and so [BR10, Theorem 19.1] gives $\mathscr{F}f\in L^{1}(\mathbb{R}^{2})$ and
@@
$\delta\coloneqq\sup_{\|\mathbf{z}\|>a}\mathopen{}\mathclose{{}\left|\mathscr{F}{f}(\mathbf{z})}\right|<1.$
@@
By properties of the Fourier transform and the product rule,
@@
$\int_{\|\mathbf{z}\|>a\sqrt{n}}\mathopen{}\mathclose{{}\left|\partial^{\alpha}\mathscr{F}{q_{t}}(\mathbf{z})}\right|\,\mathrm{d}\mathbf{z}$ $\lesssim\eta_{|\alpha|}n^{\frac{|\alpha|}{2}}\delta^{n-|\alpha|-1}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\mathscr{F}{f}\mathopen{}\mathclose{{}\left(\frac{\mathbf{z}}{\sqrt{n}}}\right)}\right|\,\mathrm{d}\mathbf{z}$ (A.3)
@@
$\lesssim\mathopen{}\mathclose{{}\left(n\beta e^{\frac{t^{2}}{2}}}\right)^{O(1)}\delta^{n-O(1)}$
$\lesssim e^{-\frac{\omega(\beta)}{4}}$
for sufficiently large $n$. Finally, we bound similar to Lemma 3.2:
@@
$\int_{\|\mathbf{z}\|>a\sqrt{n}}$ $\mathopen{}\mathclose{{}\left|\partial^{\alpha}\frac{\varphi}{6\sqrt{n}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}$ (A.4)
@@
$\lesssim n^{-\frac{1}{2}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\partial^{\alpha}H^{(k,3-k)}\varphi}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}$
$\lesssim n^{-\frac{1}{2}}\eta_{3}\int_{\mathbb{R}^{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{2}}\,\mathrm{d}\mathbf{z}$
$\lesssim e^{-\frac{\omega(\beta)}{4}}.$
Combining A.2, A.3 and A.4 proves A.1 and hence Lemma 3.4.∎
A.4 Proof of Proposition 4.1
@@
Point 1 in Theorem 2.1 can readily be seen to hold because of the explicit form of both of the fields. Point 4 also readily holds, since $F_{n}^{\prime\prime}$ is a Lipschitz function for every realization of $X_{i}$, as a sum of Lipschitz functions. We focus on showing Point 3, the proof of which can be repeated essentially verbatim to deduce Point 2.
@@
[…]
$=\frac{1}{(2\pi)^{\frac{n}{2}}}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{}\left|\int_{\mathbb{R}}e^{-\mathrm{i}(\xi_{1}g(x)+\xi_{2}g^{\prime}(x))}e^{-\frac{x^{2}}{2}}\,\mathrm{d}x}\right|^{n}\,\mathrm{d}\xi.$
Recalling that $n\to\infty$ in our regime, we can suppose $n\geqslant 4$, and for the above integral to be finite, it suffices to show that

### v3

@@ 5 Additional proofs
$\int_{0}^{\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u}^{n}e^{-\alpha u^{2}}\,\mathrm{d}u=\frac{1}{2}\Gamma\mathopen{}\mathclose{{\left(\frac{n+1}{2}}}\right)\alpha^{-\frac{n+1}{2}}.$
Moreover, for a fixed $n$, as $\epsilon\to 0$
$\int_{0}^{\infty}v^{n}e^{-(v-\epsilon)^{2}}dv\to\int_{0}^{\infty}v^{n}e^{-v^{2}}dv\quad\text{and}\quad\int_{-\infty}^{\infty}|v|^{n}e^{-(v-\epsilon)^{2}}dv\to 2\int_{0}^{\infty}v^{n}e^{-v^{2}}dv$
Proof.
For the first display, by a change of variables $v=\alpha u^{2}$, we get
$\int_{0}^{\infty}{u}^{n}e^{-\alpha u^{2}}\,\mathrm{d}u=\frac{1}{2}\alpha^{-\frac{n+1}{2}}\int_{0}^{\infty}v^{\frac{n-1}{2}}e^{-v}\,\mathrm{d}v$
and we recognize the integral as the definition of the Gamma function.
For the second display, we apply the dominated convergence theorem: clearly, the integrands converges pointwise as $\varepsilon\to 0$, and it is dominated by an integrable function that is $e2^{n}$ for $v\in[-2,2]$ and $v^{n}e^{-v^{2}/4}$ for $v\geqslant 2$ as we may take $\varepsilon\leqslant 1\leqslant|v|/2$.
5.2 Proof of Lemma˜2.2
In this section we compute the first two moments of $(G,G^{\prime})$ to prove Lemma˜2.2. Note that if $n^{c}\lesssim\beta\lesssim n^{2-c}$ for some $c>0$, and for $t\in T$, then we have $\exp\Theta(t^{2}/\beta)\to 1$. This implies that exponentials in the moments are asymptotically $e^{-{t^{2}}/{2}}$.
@@
Hence, using Lemma˜5.1 we compute
$\mathbb{E}G(t)=\int_{-\infty}^{\infty}ze^{-\frac{\beta}{2}z^{2}}\,\mathrm{d}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi(z-t)}$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}ze^{-\frac{\beta}{2}z^{2}-\frac{1}{2}(z-t)^{2}}\,\mathrm{d}z$
@@
$\mathbb{E}G^{\prime}(t)$ $=\int_{-\infty}^{\infty}(1-\beta z^{2})z^{-\frac{\beta}{2}z^{2}}\,\mathrm{d}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\varphi(z-t)}$
@@
From these computations, and the remark after Lemma˜5.1, we readily obtain the asymptotics of $\mu_{t}$ as in Lemma˜2.2 upon multiplying by $\sqrt{n}$.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}z^{2}+\frac{1}{2}(z-t)^{2}=\frac{2\beta+1}{2}u^{2}+\frac{\beta t^{2}}{2\beta+1}\quad\text{where}\quad u=z-\frac{t}{2\beta+1}.$
Hence using Lemma˜5.1 we compute
@@
It is easy to check that entries of $\Sigma_{t}$ are asymptotically the corresponding second moments. Together, we readily obtain the asymptotics of $\Sigma_{t}$ as indicated in Lemma˜2.2.∎
5.3 Proof of Lemma˜3.2
@@
In this section, we prove Lemma˜3.2 on cumulants of $Y=\Sigma_{t}^{-\frac{1}{2}}\mathopen{}\mathclose{{\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}}\right)$. To upper bound, we do not need to track the leading coefficients to ensure that they do not vanish when we combine applications of Lemma˜5.1. First, as $s$ is a constant (we only apply $s=2,3$), cumulants of order $s$ are clearly $O(\eta_{s})$. To bound $\eta_{s}$, we recall $\Sigma_{t}^{-1}$ from the proof of Lemma˜2.3 and apply Hölder’s inequality:
$\eta_{s}$ $=\mathbb{E}\mathopen{}\mathclose{{\left[\|Y\|^{s}}}\right]$
@@
$\leqslant\mathbb{E}\mathopen{}\mathclose{{\left\|\mathopen{}\mathclose{{\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}}\right)^{\intercal}\Sigma_{t}^{-1}\mathopen{}\mathclose{{\left(G-\mathbb{E}G,G^{\prime}-\mathbb{E}G^{\prime}}}\right)}}\right\|^{\frac{s}{2}}$
@@
$\lesssim\beta^{\frac{s}{4}}e^{\frac{st^{2}}{4}}\mathbb{E}\mathopen{}\mathclose{{\left|\beta(G-\mathbb{E}G)^{2}+2t(G-\mathbb{E}G)(G^{\prime}-\mathbb{E}G^{\prime})+(G^{\prime}-\mathbb{E}G^{\prime})^{2}}}\right|^{\frac{s}{2}}$
@@
$\lesssim\beta^{\frac{s}{4}}e^{\frac{st^{2}}{4}}\mathopen{}\mathclose{{\left(\beta^{\frac{s}{2}}\mathbb{E}|G|^{s}+\mathbb{E}|G^{\prime}|^{s}}}\right)$
@@
$\lesssim\beta^{\frac{s}{4}}e^{\frac{(s-2)t^{2}}{4}}\int_{0}^{\infty}h_{s}\mathopen{}\mathclose{{\left(|z|}}\right)e^{-\frac{3\beta+1}{2}\mathopen{}\mathclose{{\left(z-\frac{t}{3\beta+1}}}\right)^{2}}\,\mathrm{d}z,$
@@
where $h_{s}(u)\coloneqq\beta^{\frac{s}{2}}u^{s}+(1+\beta u^{2})^{s}$. Note that the shift $t/(3\beta+1)\ll((3\beta+1)/2)^{-1/2}$, so by linearity of integration, we may apply Lemma˜5.1 to bound the integral of each monomial $|z|^{\ell}$ by $O(\beta^{-(\ell+1)/2})$. By monotonicity of $h$ on $\mathbb{R}_{\geqslant 0}$ and since $t\in T$, for some constant $C>0$,
$\eta_{s}\lesssim\beta^{\frac{s-2}{4}}e^{\frac{(s-2)t^{2}}{4}}h\mathopen{}\mathclose{{\left(C\beta^{-\frac{1}{2}}}}\right)\lesssim\beta^{\frac{s-2}{4}}e^{\frac{(s-2)t^{2}}{4}}$
@@
upon noting $u\mapsto h(u/\sqrt{\beta})$ has constant coefficients. This proves Lemma˜3.2.
5.4 Proof of Lemma˜3.5
@@
Fix $n,\beta$ sufficiently large as well as $t$. Define for a multi-index $\alpha$ that $h(\mathbf{x})=\mathbf{x}^{\alpha}g_{s}(\mathbf{x})$. Then, for $s=3$
@@
$\mathscr{F}h(\mathbf{z})=\partial^{\alpha}\mathopen{}\mathclose{{\left(\mathscr{F}(q_{t})-2\pi\varphi-\sum_{k=0}^{3}\frac{\varphi}{k!(3-k)!\sqrt{n}}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}}}\right)(\mathbf{z}),$
and for $s=2$ we have the same statement with the last summand omitted. Note that we use
$\mathscr{F}h(\mathbf{z})\coloneqq\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\langle\mathbf{x},\mathbf{z}\rangle}h(\mathbf{x})\,\mathrm{d}\mathbf{x}$
@@
to denote the Fourier transform of $h$. We also omit the dependence of $h$ on $s$ and $\alpha$ for brevity. By Fourier inversion, it suffices to show that for any multi-index $\alpha$ with order $|\alpha|\leqslant s$ that
@@
$\mathopen{}\mathclose{{\left|h(\mathbf{x})}}\right|=\mathopen{}\mathclose{{\left|\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}e^{-\mathrm{i}\langle\mathbf{z},\mathbf{x}\rangle}\mathscr{F}{h}(\mathbf{z})\,\mathrm{d}\mathbf{z}}}\right|\lesssim\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{\left|\mathscr{F}{h}(\mathbf{z})}}\right|\,\mathrm{d}\mathbf{z}\lesssim n^{-\frac{s-1}{2}}\eta_{s+1}$ (5.1)
@@
We apply [bhattacharya2010normal, Theorem 9.10]—which is not asymptotic and has explicit constants in $s$ only—so we may use it even though $q_{t}$ depends on $\beta$ to obtain that
@@
$\mathopen{}\mathclose{{\left|\mathscr{F}{h}(\mathbf{z})}}\right|\lesssim n^{-\frac{s-1}{2}}\eta_{s+1}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{4}}$ (5.2)
@@
provided $\|\mathbf{z}\|\leqslant a\sqrt{n}$ for some $a\asymp\eta_{s+1}^{-\frac{1}{s-1}}$. By Lemma˜3.2, we have that
@@
$\int_{\|\mathbf{z}\|\leqslant a\sqrt{n}}\mathopen{}\mathclose{{\left|\mathscr{F}{h}(\mathbf{z})}}\right|\,\mathrm{d}\mathbf{z}\lesssim n^{-\frac{s-1}{2}}\eta_{s+1}\int_{\mathbb{R}^{2}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{4}}\,\mathrm{d}\mathbf{z}\lesssim n^{-\frac{s-1}{2}}\eta_{s+1}.$ (5.3)
Recall that $q_{t}$ is the density of $n^{-\frac{1}{2}}\sum_{i=1}^{n}Y_{i}$. Let $f$ denote the density of $W\coloneqq\sqrt{5}(Y_{1}+\dots+Y_{5})$ which exists and is bounded by Proposition˜4.1. Hence, $\mathscr{F}f\in L^{1}(\mathbb{R}^{2})$ and
@@
$\varepsilon\coloneqq\sup_{\|\mathbf{z}\|>a}\mathopen{}\mathclose{{\left|\mathscr{F}{f}(\mathbf{z})}}\right|<1.$
@@
Now, $q_{t}$ is the density of $\sqrt{n/5}$ times the sum of $n/5$ many i.i.d. copies of $W$, so by properties of the Fourier transform and the product rule,
@@
$\int_{\|\mathbf{z}\|>a\sqrt{n}}\mathopen{}\mathclose{{\left|\partial^{\alpha}\mathscr{F}{q_{t}}(\mathbf{z})}}\right|\,\mathrm{d}\mathbf{z}$ $\lesssim\eta_{|\alpha|}n^{\frac{|\alpha|}{2}}\varepsilon^{n/5-|\alpha|-1}\int_{\mathbb{R}^{2}}\mathopen{}\mathclose{{\left|\mathscr{F}{f}\mathopen{}\mathclose{{\left(\frac{\mathbf{z}}{\sqrt{n}}}}\right)}}\right|\,\mathrm{d}\mathbf{z}$ (5.4)
@@
$\lesssim\mathopen{}\mathclose{{\left(n\beta e^{\frac{t^{2}}{2}}}}\right)^{O(1)}\varepsilon^{n/5-s-1}$
$\ll n^{-\frac{s-1}{2}}\eta_{s+1}$
for sufficiently large $n$. Finally, we bound similar to Lemma˜3.3:
@@
$\int_{\|\mathbf{z}\|>a\sqrt{n}}\mathopen{}\mathclose{{\left|\partial^{\alpha}}}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}\lesssim\int_{{\|\mathbf{z}\|>a\sqrt{n}}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{2}}\,\mathrm{d}\mathbf{z}\ll n^{-\frac{s-1}{2}}\eta_{s+1}$ (5.5)
and for the $s=3$ we also have the additional term
@@
$\int_{\|\mathbf{z}\|>a\sqrt{n}}$ $\mathopen{}\mathclose{{\left|\partial^{\alpha}\sum_{k=0}^{3}\frac{\varphi}{k!(3-k)!\sqrt{n}}\kappa_{t}^{(k,3-k)}H^{(k,3-k)}}}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}$ (5.6)
$\lesssim n^{-\frac{1}{2}}\sum_{k=0}^{3}\kappa_{t}^{(k,3-k)}\int_{{\|\mathbf{z}\|>a\sqrt{n}}}\mathopen{}\mathclose{{\left|\partial^{\alpha}H^{(k,3-k)}\varphi}}\right|(\mathbf{z})\,\mathrm{d}\mathbf{z}$
$\lesssim n^{-\frac{1}{2}}\eta_{3}\int_{{\|\mathbf{z}\|>a\sqrt{n}}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{2}}\,\mathrm{d}\mathbf{z}$
$\ll n^{-\frac{s-1}{2}}\eta_{s+1}$
Now, in the last step of both ˜5.5 and 5.6, we use that the standard Gaussian integral outside the ball at the origin converges to zero exponentially quickly as radius $a\sqrt{n}\to\infty$ by ˜3.6, so in particular
$\int_{{\|\mathbf{z}\|>a\sqrt{n}}}\|\mathbf{z}\|^{O(1)}e^{-\frac{\|\mathbf{z}\|^{2}}{2}}\,\mathrm{d}\mathbf{z}\ll(a\sqrt{n})^{-(s-1)}\asymp n^{-\frac{s-1}{2}}\eta_{s+1}$
Combining ˜5.3, 5.4, 5.5 and 5.6 proves ˜5.1 and hence Lemma˜3.5 for both $s=2$ and $s=3$ cases. ∎
5.5 Proof of Proposition˜4.1
@@
Point 1 in Theorem˜2.1 can readily be seen to hold because of the explicit form of both of the fields. Point 4 also readily holds, since $F_{n}^{\prime\prime}$ is a Lipschitz function for every realization of $X_{i}$, as a sum of Lipschitz functions. We focus on showing Point 3, the proof of which can be repeated essentially verbatim to deduce Point 2.
@@
We write $\xi=(\xi_{1},\xi_{2})$ with $\|\xi\|=\rho$, and $\omega=(\cos\theta,\sin\theta)$ so that $\xi=\rho\omega$. We then have
[…]
with phase $\phi_{\theta}(x)=\cos\theta g(x)+\sin\theta g^{\prime}(x)$. We will show the uniform bound
$\mathopen{}\mathclose{{\left|\mathscr{F}\nu_{t}(\xi)}}\right|\lesssim\frac{1}{\sqrt{1+\|\xi\|}},$ (5.7)
for all $\xi\in\mathbb{R}^{2}$, where the implicit constant depends only on $\beta$. To this end, we compute
$g^{\prime}(x)$ $=(1-\beta x^{2})e^{-\beta\frac{x^{2}}{2}},$
$g^{\prime\prime}(x)$ $=\beta x(\beta x^{2}-3)e^{-\beta\frac{x^{2}}{2}},$
$g^{\prime\prime\prime}(x)$ $=\beta(-\beta^{2}x^{4}+6\beta x^{2}-3)e^{-\beta\frac{x^{2}}{2}}.$
Set $\psi(x)=(g(x),g^{\prime}(x))$. We have
$\det(\psi^{\prime}(x),\psi^{\prime\prime}(x))=g^{\prime}(x)g^{\prime\prime\prime}(x)-(g^{\prime\prime}(x))^{2}=-\beta(\beta^{2}x^{4}+3)e^{-\beta x^{2}}<0$
for all $x.$ In particular the determinant never vanishes. Observe that $\phi_{\theta}^{\prime}(x)=\cos\theta g^{\prime}(x)+\sin\theta g^{\prime\prime}(x)$, and any stationary point $x_{0}$ of $\phi_{\theta}$ satisfies $g^{\prime}(x_{0})=0$. If also $\phi^{\prime\prime}(x_{0})=\cos\theta g^{\prime\prime}(x_{0})+\sin\theta g^{\prime\prime\prime}(x_{0})=0$, then $(\cos\theta,\sin\theta)$ would be a nontrivial vector orthogonal to both $(g^{\prime}(x_{0}),g^{\prime\prime}(x_{0}))$ and $(g^{\prime\prime}(x_{0}),g^{\prime\prime\prime}(x_{0}))$, forcing $g^{\prime}(x_{0})g^{\prime\prime\prime}(x_{0})-(g^{\prime\prime}(x_{0}))^{2}=0$, a contradiction. Hence all stationary points of $\phi_{\theta}$ are non-degenerate, uniformly in $\theta$.
Fix $R>0$ and a smooth cutoff function $\chi\in C_{c}^{\infty}(\mathbb{R})$ with $\chi\equiv 1$ on [1,1][-1,1], supported within [2,2][-2,2]. Then set $\chi_{R}(x)=\chi(x/R)$. We write
$\mathscr{F}\nu_{t}(\xi)$ $=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}e^{-\mathrm{i}\rho\phi_{\theta}(x)}\chi_{R}(x)e^{-\frac{x^{2}}{2}}\,\mathrm{d}x+\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}e^{-\mathrm{i}\rho\phi_{\theta}(x)}(1-\chi_{R}(x))e^{-\frac{x^{2}}{2}}\,\mathrm{d}x$
$=I_{1}+I_{2}.$
@@
$|I_{2}|\leqslant\int_{|x|\geqslant R}e^{-\frac{x^{2}}{2}}\leqslant\frac{1}{\sqrt{2\pi}}\frac{2}{R}e^{-\frac{R^{2}}{2}},$
which, as $R$ is fixed, is smaller than $(1+\rho)^{-\frac{1}{2}}$ whenever $\rho$ is large enough, uniformly in $\theta$. Concerning $I_{1}$, because $\phi^{\prime}_{\theta}$ is analytic it only has finitely many zeros $x_{1}(\theta),\ldots,x_{m}(\theta)$ on [2R,2R][-2R,2R] (with $m\leqslant 3$). They are all non-degenerate: $\phi^{\prime\prime}_{\theta}(x_{j}(\theta))\neq 0$. Therefore there exist disjoint intervals $J_{j}(\theta)\subset[-2R,2R]$ around each $x_{j}(\theta)$ of radius $\delta>0$, and some $c_{*}=c_{*}(R,\beta,\delta)>0$ such that $|\phi^{\prime\prime}_{\theta}(x)|\geqslant c_{*}$ for all $x$ in the union of the intervals $J_{j}(\theta)$ and all $\theta\in\mathbb{S}^{1}$. On each $J_{j}(\theta)$, we may apply the method of stationary phase to find